### v2

@@ 2 Basic AI Techniques for Embodied AI
EmAI is gaining momentum thanks to advancements across multiple fields, particularly the breakthroughs in AI. To ultimately replicate human-like behavior in real-world contexts [126], a comprehensive EmAI “brain” should encompass multiple modules to conduct perception, action control, decision-making, and memory. Similar to the human brain, which consists of several specialized but interconnected functional regions (see Figure 2(a)), these integrated capabilities enable EmAI systems to interact with and adapt to complex real-world environments [127, 128, 129, 124], as illustrated in Figure 2(b). Here, we outline key approaches that support these functions, categorized into embodied perception, low-level actuation, high-level planning, and memory processing, along with their detailed breakdowns, as shown in Figure 3. We will also summarize the major achievements in this section.
@@ 2.1 Embodied Perception
Perception is a core mechanism by which EmAI systems interpret sensory data from their environment. This process involves handling high-dimensional, multimodal, and often noisy inputs from sensors such as cameras, microphones, and tactile devices. This section classifies embodied perception from three key aspects: sensory perception, cross-modal perception, and interactive perception. Sensory perception forms the basis for other system functionalities and integrates directly with existing single-modality foundation models [130, 131, 132, 133, 134, 135, 136, 54, 55, 80]. To achieve a richer understanding of the environment, multimodal AI algorithms [137, 138, 139, 140] enable cross-modal information integration, aligning with the inherently multimodal nature of the real world. The integration of multimodal data from different devices allows robots to combine sensory inputs such as vision, touch, and speech to make more informed decisions [141]. Interactive perception, which further learns object affordances, represents a pivotal step in bridging perception and action, highlighting its significance as the next frontier in EmAI perception. We will delve into cross-modal and interactive perception in greater detail in the following sections.
@@ 2.1.1 Cross-modal Perception
Cross-modal perception integrates information from multiple modalities to achieve a holistic understanding. To efficiently aggregate and align multimodal information, current pre-trained models establish foundational cross-modal representations, enabling downstream multimodal tasks that can be reformulated into challenges such as Visual Language Navigation (VLN) and Embodied Visual Question Answering (VQA), etc. Recent studies [142, 143, 144, 145, 146, 147, 148, 149] predominantly adopt three primary architectural paradigms to achieve effective cross-modal perception:
@@
Deeper cross-modality interaction is often achieved by employing a shared encoder to learn comprehensive cross-modal representations. These shared encoders typically leverage multi-layer Transformers that encode multimodal inputs through representation fusion techniques such as cross-attention mechanisms [162, 163, 149, 164] or feature-wise linear modulation (FiLM) [145, 165, 146]. Representative shared encoder frameworks include ViLBERT [166], LXMERT [167], and UNITER [168]. By using deep fusion encoders, this architecture learns more generic cross-modal representations and achieves superior pre-training efficacy for visual reasoning tasks [169, 170, 171, 172] and few-shot tasks [173, 174, 175, 176, 177]. However, it introduces quadratic (for two modalities) or cubic (for three modalities) time complexity, as it requires interaction between all possible modality pairs. This results in significantly slower inference speeds, limiting its practical applicability. Moreover, bridging the cross-modality gap with a single parameterized model remains challenging, further complicating its deployment in real-world scenarios.
Therefore, a combination of the aforementioned approaches, referred to as the combination architecture, has also been proposed. Typically, these methods employ separate encoders for modality-specific feature extraction alongside a shared encoder for joint feature learning. Furthermore, techniques aimed at reducing computational overhead, such as Mixture-of-Modality-Experts (MoME) [157, 143, 178, 179], Mixture-of-Prompt-Experts (MoPE) [180], effective self-attention [181, 182], and Multi-directional Adapter [147] utilize sparse routing or modular expert networks to efficiently manage modality-specific information while minimizing computational costs.
@@ 2.1.2 Interactive Perception
Interactive perception involves physical actions—such as manipulating objects, changing viewpoints, or probing the environment—to resolve ambiguities, learn object properties, and refine multimodal representations [183, 184, 185]. By leveraging exploratory actions, it allows EmAI systems to enhance or extend their perception abilities in object recognition [185], scene understanding [186], or manipulation in dynamic and unstructured environments [183], etc. In robotic manipulation, interactive perception gathers data through exploration to identify potential actionable regions of objects and understand their functional possibilities. This process, a.k.a. affordance learning, further benefits EmAI systems by guiding and optimizing their future interactions.
With affordance learning, action plans can be nominated through two key approaches: supervised learning from human demonstrations [187, 188] and reinforcement learning from robotic trial-and-error interactions [189, 190]. Pioneering approaches like Where2Act [191] enable robots to identify the most effective interaction strategies for various parts of an object, and Where2Explore [192] generalize affordance knowledge to similar object parts, enabling robots to adapt to unseen objects with limited prior experience. Additionally, by integrating reinforcement learning, RLAfford [193] facilitates end-to-end affordance learning, enabling robots to adapt seamlessly to a wide range of manipulation tasks.
@@ 2.2 Low-level Actuation
Low-level actuation is a fundamental component of EmAI systems that leverages various action control policies to determine real-time motor control based on perceptions [194]. In this section, we explore the process of low-level actuation by dividing it into two core phases: control policy representation and control policy learning. The phase of policy representation showcases the framework for encoding robotic behaviors, ensuring that policies are expressive enough to capture intricate actions while remaining computationally efficient and adaptable to diverse scenarios. Building on this foundation, the policy learning phase focuses on how robots select and optimize these behaviors through advanced algorithms including reinforcement learning, imitation learning, and hybrid strategies. Together, these two phases form a cohesive framework that equips robots with the ability to act autonomously and achieve predefined objectives. Table 1 summarizes representative low-level control policies.
@@ 2.2.1 Policy Representation
Three types of policy representations were widely used in current EmAI systems. I) Explicit policies [195, 196] is a common approach in behavior cloning [236] that maps directly from observations to actions. It can be viewed as a regression task designed to learn a mapping function, typically implemented via neural networks, to reproduce the behavior of the demonstrators. According to the output actions, deterministic policies [197, 237, 198] provide fixed options, while stochastic policies [199, 238] encourage exploration by introducing randomness or initializing with high entropy. While straightforward to implement, explicit policies struggle with complex tasks due to insufficient expressive capability. II) Implicit policies [204, 239] such as Energy-Based Models [201, 202, 203], they indirectly represent the policy through an energy function that defines the preference distribution in the state-action space, making them more suitable for multimodal distribution tasks. III) Diffusion policies, inspired by Denoising Diffusion Probabilistic Models (DDPMs) [240], create a policy with the conditional generative model [212] and generate actions by refining noise, offering multimodal, expressive representations with applications in offline reinforcement learning [241] and vision-based manipulation [242]. While promising, diffusion policies require further optimization to improve sampling efficiency and inference speed in complex action spaces.
@@ 2.2.2 Policy Learning
Reinforcement Learning (RL) methods aim to optimize policies through trial-and-error interactions with the environment, guided by a scalar reward signal that encodes task performance. Recent RL algorithms can be broadly categorized into value-based, policy-based, and actor-critic approaches. Value-based methods, such as Deep Q-Networks (DQN) [247], estimate action values using temporal difference learning, enabling discrete action selection based on the greedy policy. In contrast, policy-based methods, like Proximal Policy Optimization (PPO) [248] and Trust Region Policy Optimization (TRPO) [249], directly optimize the policy in continuous action spaces by minimizing divergence between successive policies to ensure stable learning. Actor-critic methods [250] combine these paradigms, where an actor (policy) generates actions, and a critic (value function) evaluates them, providing gradients to optimize the policy iteratively. However, RL algorithms often suffer from challenges such as sampling inefficiency, high exploration costs, and instability, which hinder their practical applications in complex environments.
Imitation learning (IL) provides an alternative by directly learning from expert demonstrations. These algorithms bypass the need for explicit reward functions, training policies to replicate expert behaviors using supervised learning techniques. Methods such as behavioral cloning [204, 251] minimize the discrepancy between predicted and expert actions, while inverse reinforcement learning [252, 253] aims to infer the underlying reward function from demonstrations. Although IL accelerates learning by reducing exploration, it struggles with covariate shift, where small errors during policy execution compound over time, limiting generalization in novel scenarios [254].
A hybrid approach that combines RL and IL leverages the strengths of both. Typically, policies are initialized with expert demonstrations to accelerate training using IL, and then refined by RL to adapt to dynamic environments [255, 254, 256]. There are also approaches such as Guided Policy Search [233] or Reward Shaping [234], which integrate demonstration data into the RL process, enhancing both data efficiency and generalization. These combinations bridge the gap between the efficiency of the IL and the adaptability of the RL, improving the low-level actuation capability to execute actions in real-world scenarios.
@@ 2.3 High-level Planning
Low-level actuation can only meet the needs of simple, reactive tasks, but it struggles to handle the complexity of planning long-horizon tasks with multiple sub-tasks. To address this limitation, high-level planning algorithms have been developed. Classic planning algorithms, such as A* algorithm [257, 258, 259, 260], Dijkstra’s algorithm [261, 262, 263], and the probabilistic roadmap (PRM) approaches [264, 265, 266, 267], serve as foundational methods in this domain. Despite their significant influence and effectiveness in structured environments, these algorithms encounter substantial challenges in real-world scenarios, particularly in high-dimensional state spaces and under conditions of partial observability. Recent research employs LLMs as high-level planners in embodied systems, bridging cognitive reasoning and physical task execution by translating abstract instructions into actionable robotic tasks [268, 65]. Table 2 summarizes representative approaches of different high-level planning methods.
@@
LLM-based task planners typically break down high-level goals into a sequence of executable subtasks [269, 270, 271]. Typically, there are two main paradigms: code-based planners and language-based planners, as shown in the Figure 4. Code-based planners [272, 273, 274] operate by selecting from a pre-defined set of modular skills or functions, invoking them via APIs to execute tasks step by step. They excel in highly controlled workflows requiring safety and reliability, where predefined APIs or modules can ensure predictable outcomes. Representative code-based planners such as the DEPS framework [275] further emphasize “self-explanation” to better exploit the capabilities of LLMs, where an LLM generates plans, explains failures, and uses environmental feedback to aid in re-planning. In addition, by transforming observation sequences into 3D scene graphs using visual language models, ConceptGraphs [274] helps the LLM reason about spatial and semantic relationships in task planning. These methods benefit from the deterministic nature of pre-programmed functions, reducing ambiguity in execution. However, their dependency on predefined skills makes them less adaptable to unexpected changes or tasks outside the pre-programmed domain.
@@
Language-based planners [276, 277, 278] show better flexibility without using predefined functions. In open-ended domains like creative problem-solving, customer support, or planning in uncertain environments, language-based planners can use contextual cues to refine and adjust their instructions in real-time. Additionally, these methods dynamically integrate diverse feedback—ranging from success indicators to human inputs—without necessitating additional training for the LLMs. These adaptive feedback and re-planning mechanisms [279, 275, 280] enable systems to recover from unexpected states, offering more flexibility to novel environments. This adaptability, however, can sometimes come at the cost of consistency or precision [123]. Moreover, by integrating MLLMs’ capabilities, EmAI systems like Socratic models [281] effectively translate non-language inputs into unified language descriptions through multimodal informed prompting. This approach not only streamlines information exchange across different modalities through MLLMs but also enhances robot perception and planning tasks [282, 123, 283].
@@ 2.5 Synergistic Integration
The aforementioned four key functionalities are often initially developed independently but must be effectively integrated to construct a comprehensive EmAI system. Integration approaches such as MemoRAG [336] and Reflexion [337] enhance high-level planning processes by retrieving relevant information [338, 339, 340] or summarizing past experiences [341, 342] from memory modules. These methods improve adaptability to dynamic environments and enable more rational decision-making. Closed-loop approaches, including RoboGolf [343], LyRN [344], and AlphaBlock [345], integrate perception and actuation modules, leveraging feedback to refine observations and dynamically update control signals. This integration facilitates precise action adjustments and supports effective multi-step planning. Besides, active and interactive perception systems [185, 44, 346] take this a step further by engaging in real-time interactions to explore object properties, update environmental contexts, and refine decisions based on immediate outcomes. The resulting actions and observations are stored in the EmAI memory, which can be used to construct multimodal knowledge graphs integrating physical properties, concepts, affordances, and intentions for future use [347]. The modern AI alignment approaches that combine modules [348, 349] and foundational models that encompass all functionalities of these modules [350, 351] are also being widely researched as promising areas [352, 353]. However, there is still a lack of a highly compatible, efficient, and effective unified architecture capable of integrating various developed modules. Achieving alignment and seamless integration among these modules, while minimizing development (e.g., fine-tuning), remains an open challenge.
@@ 3.1.1 Pre-Intervention Stage
Virtual triage nurse. In modern smart healthcare systems, EmAI-based virtual triage nurses replace human nurses and play a pivotal role in streamlining patient management by directing individuals to the most appropriate clinical departments. These EmAI systems analyze patient-reported symptoms and conduct department ranking based on symptom descriptions [359, 360, 361]. Beyond symptom-based sorting, advanced triage systems integrate wearable health data and EHR to provide a holistic assessment of conditions [362, 363, 364]. In emergency care settings, these systems can even predict patient outcomes and recommend intervention pathways, significantly shortening response times during critical situations [365, 366, 367, 368]. As healthcare systems become more interconnected, intelligent triage systems are increasingly serving as an efficient tool of future infrastructure [364, 369], enabling seamless coordination between primary care, specialist consultations, and hospital admissions.
@@
Interactive medical consultant. Recently, some LLM-based chatbots, such as DISC-MedLLM [370] and HealAI [371], are used to provide instant, reliable, and context-specific responses to medical inquiries, helping patients better understand their symptoms, treatment options, or follow-up care [372, 372, 373]. They can also explain medical conditions, offer personalized recommendations [374, 375], and explain radiology reports [376]. By bridging the gap between patients and complex medical knowledge, interactive medical chatbots not only empower individuals to make informed decisions but also reduce the workload of healthcare professionals [377, 378]. With advancements in LLM reasoning [379], these systems are becoming an essential component of patient-centered smart healthcare solutions. Beyond answering medical questions, these chatbots can also guide patients through administrative processes, such as booking appointments [380, 381] or managing prescriptions [382, 383], thereby easing the burden on healthcare staff [384, 385]. With their ability to operate around the clock, these systems foster greater accessibility and trust in the medical process.
@@
Imaging analyst. Another representative area is robot-assisted medical imaging, which not only improves the accuracy of diagnostic processes but also expands the capabilities of medical imaging in complex anatomical assessments. There have been many studies focused on medical image analysis using AI techniques [386, 387, 388], but to support robotic surgery and preoperative robotic diagnosis requires additional requirements including real-time processing [389], 3D spatial understanding [390], and safety compliance [391] beyond high precision. Among medical imaging technologies including ultrasonography, radiology, and endoscopy, EmAI enhances their capabilities in different ways.
@@
For ultrasonography, as it is portable, real-time, non-invasive, and relying on the synchronization of diagnostic and procedural operations, EmAI can serve as a remote assistant to help clinicians conduct remote ultrasound diagnosis [105] and protect themselves from risks against epidemics [392]. For radiology, EmAI can also help clinicians in lesion localization [77, 393, 394], surgical planning [395, 396, 397], intraoperative navigation [398, 399] with a better understanding of the raw imaging and attached report from X-Ray, Computed Tomography (CT), and Magnetic Resonance Imaging (MRI). For endoscopy, EmAI can process endoscopic video feeds in real-time to identify abnormalities [400, 401], such as polyps or early-stage tumors, reduce operator fatigue, enable precise navigation via path optimization [402, 403], and assist in procedural tasks like polypectomy with adaptive, precision-controlled robotic movements [404, 405].
@@
    Endoscopic navigator. The integration of EmAI in endoscopic procedures has significantly enhanced the precision and efficiency of minimally invasive diagnostics and treatments by providing real-time guidance for device positioning and lesion targeting [426, 427]. For example, AI-powered 3D mapping techniques reconstruct anatomical structures in real time, enabling clinicians to navigate complex regions such as the gastrointestinal tract and respiratory pathways with greater accuracy [428]. Robotic-assisted endoscopy further leverages automated control systems that adapt dynamically to patient-specific anatomy and procedural demands [429, 430], thereby improving precision in tasks such as biopsy, polyp removal, and other targeted interventions. Furthermore, these systems incorporate predictive analytics to refine surgical pathways, reduce tissue trauma, and streamline procedural workflows, enhancing safety and operator confidence in endoscopic interventions [431, 432].
In summary, fully autonomous ultrasound system [110], human-centric radiology assistant [433], and flexible robotic endoscope system [434] capable of navigating and adjusting diagnostic examinations with minimal human supervision have been developed, achieving a higher level of intelligence and autonomy. Compared to human-led examinations, they can optimize workflows, improve patient care quality, and ensure a safer environment for all involved in the surgical process. Besides, these systems can also function as a module within a more comprehensive EmAI system such as surgical robots, showcasing opportunities to expand their applications.

### v3

@@ 2 Basic AI Techniques for Embodied AI
EmAI is gaining momentum thanks to advancements across multiple fields, particularly recent breakthroughs in AI. To ultimately replicate human-like behavior in real-world contexts [126], a comprehensive EmAI “brain” should encompass multiple modules that conduct perception, action control, decision-making, and memory. Similar to the human brain, which consists of several specialized but interconnected functional regions (see Figure 2(a)), these integrated capabilities enable EmAI systems to interact with and adapt to complex real-world environments [127, 128, 129, 124], as illustrated in Figure 2(b). Here, we outline key approaches that support these functions, categorized into embodied perception, low-level actuation, high-level planning, and memory processing, along with their breakdowns in detail, as shown in Figure 3. We will also summarize the major achievements in these areas in this section.
@@ 2.1 Embodied Perception
Perception is a core mechanism by which EmAI systems interpret sensory data from their environment. This process involves handling high-dimensional, multimodal, and often noisy inputs from sensors such as cameras, microphones, and tactile devices. This section divides embodied perception from three key aspects: sensory perception, cross-modal perception, and interactive perception. Sensory perception forms the basis for other system functionalities and integrates directly with existing single-modality foundation models [130, 131, 132, 133, 134, 135, 136, 53, 54, 80]. To achieve a richer understanding of the environment, multimodal AI algorithms [137, 138, 139, 140] enable the integration of cross-modal information, aligning with the inherently multimodal nature of the real world. The perception of multimodal data from different devices allows robots to combine sensory inputs such as vision, touch, and speech to make more informed decisions [141]. Interactive perception, which further explores object affordances, represents a pivotal step in bridging perception and action, making it a significant area for future development in EmAI perception. We will delve into cross-modal and interactive perception in greater detail in the following sections.
@@ 2.1.1 Cross-modal Perception
Cross-modal perception integrates information from multiple modalities to achieve a holistic understanding. To efficiently aggregate and align multimodal information, current pre-trained models establish foundational cross-modal representations, enabling downstream multimodal tasks such as Visual Language Navigation (VLN) and Embodied Visual Question Answering (VQA), etc. Recent studies [142, 143, 144, 145, 146, 147, 148, 149] predominantly adopt three primary architectural paradigms to achieve effective cross-modal perception:
@@
Deeper cross-modality interaction is often achieved by employing a shared encoder to learn comprehensive cross-modal representations. These shared encoders typically leverage multi-layer Transformers that encode multimodal inputs through representation fusion techniques such as cross-attention mechanisms [162, 163, 149, 164] or feature-wise linear modulation (FiLM) [145, 165, 146]. Representative shared encoder frameworks include ViLBERT [166], LXMERT [167], and UNITER [168]. By using deep fusion encoders, this architecture learns more generic cross-modal representations and achieves superior pre-training efficacy for visual reasoning tasks [169, 170, 171, 172] and few-shot tasks [173, 174, 175, 176, 177]. However, it introduces quadratic (for two modalities) or cubic (for three modalities) time complexity, as it requires interaction between all possible modality pairs. This results in significantly slower inference speeds, which limits its practical applicability. Moreover, bridging the cross-modality gap with a single parameterized model remains challenging, further obstructing its deployment in real-world scenarios.
Therefore, a combination of the aforementioned approaches, referred to as the combination architecture, has also been proposed. Typically, these methods employ separate encoders for modality-specific feature extraction alongside a shared encoder for joint feature learning. Furthermore, techniques aimed at reducing computational overhead, such as Mixture-of-Modality-Experts (MoME) [157, 143, 178, 179], Mixture-of-Prompt-Experts (MoPE) [180], effective self-attention [181, 182], and Multi-directional Adapter [147], use sparse routing or modular expert networks to efficiently manage modality-specific information while minimizing computational costs.
@@ 2.1.2 Interactive Perception
Interactive perception involves physical actions—such as manipulating objects, changing viewpoints, or probing the environment—to resolve ambiguities, learn object properties, and refine multimodal representations [183, 184, 185]. By leveraging exploratory actions, it allows EmAI systems to enhance or extend their perception abilities in object recognition [185], scene understanding [186], or manipulation in dynamic and unstructured environments [183], among others. In robotic manipulation, interactive perception gathers data through exploration to identify potential actionable regions of objects and understand their functional possibilities. This process, a.k.a. affordance learning, further benefits EmAI systems by guiding and optimizing their future interactions.
With affordance learning, action plans can be generated through two key approaches: supervised learning from human demonstrations [187, 188] and reinforcement learning from robotic trial-and-error interactions [189, 190]. Pioneering approaches like Where2Act [191] enable robots to identify the most effective interaction strategies for various parts of an object, and Where2Explore [192] generalize affordance knowledge to similar object parts, enabling robots to adapt to unseen objects with limited prior experience. Additionally, by integrating reinforcement learning, RLAfford [193] facilitates end-to-end affordance learning, enabling robots to seamlessly adapt to a wide range of manipulation tasks.
@@ 2.2 Low-level Actuation
Low-level actuation is a fundamental component of EmAI systems that leverages various action control policies to determine real-time motor control based on perceptions [194]. In this section, we explore the process of low-level actuation by dividing it into two core phases: control policy representation and control policy learning. The phase of policy representation presents the framework for encoding robotic behaviors, ensuring that policies are expressive enough to capture intricate actions while remaining computationally efficient and adaptable to diverse scenarios. Building on this foundation, the policy learning phase focuses on how robots select and optimize these behaviors through advanced algorithms including reinforcement learning, imitation learning, and hybrid strategies. Together, these two phases form a cohesive framework that equips robots with the ability to act autonomously and achieve predefined objectives. Table 1 summarizes representative low-level control policies.
@@ 2.2.1 Policy Representation
Three types of policy representations are widely used in current EmAI systems. I) Explicit policies [195, 196] is a common approach in behavior cloning [236] that maps directly from observations to actions. This approach can be viewed as a regression task designed to learn a mapping function, typically implemented via neural networks, to reproduce the behavior of the demonstrators. According to the output actions, deterministic policies [197, 237, 198] provide fixed options, while stochastic policies [199, 238] encourage exploration by introducing randomness or initializing with high entropy. While straightforward to implement, explicit policies struggle with complex tasks due to insufficient expressive capability. II) Implicit policies [204, 239], such as Energy-Based Models [201, 202, 203], indirectly represent the policy through an energy function that defines the preference distribution in the state-action space, making them more suitable for multimodal distribution tasks. III) Diffusion policies, inspired by Denoising Diffusion Probabilistic Models (DDPMs) [240], create a policy using the conditional generative model [212] and generate actions by refining noise, offering multimodal, expressive representations with applications in offline reinforcement learning [241] and vision-based manipulation [242]. While promising, diffusion policies require further optimization to improve sampling efficiency and inference speed in complex action spaces.
@@ 2.2.2 Policy Learning
Reinforcement Learning (RL) methods aim to optimize policies through trial-and-error interactions with the environment, guided by a scalar reward signal that encodes task performance. Recent RL algorithms can be broadly categorized into value-based, policy-based, and actor-critic approaches. Value-based methods, such as Deep Q-Networks (DQN) [247], estimate action values using temporal difference learning, enabling discrete action selection based on the greedy policy. In contrast, policy-based methods, like Proximal Policy Optimization (PPO) [248] and Trust Region Policy Optimization (TRPO) [249], directly optimize the policy in continuous action spaces by minimizing divergence between successive policies to ensure stable learning. Actor-critic methods [250] combine these paradigms, where an actor (policy) generates actions, and a critic (value function) evaluates them, providing gradients to optimize the policy iteratively. However, RL algorithms often suffer from challenges such as sampling inefficiency, high exploration costs, and instability, which limit their practical applications in complex environments.
Imitation learning (IL) provides an alternative by directly learning from expert demonstrations. These algorithms bypass the need for explicit reward functions, training policies to replicate expert behaviors using supervised learning techniques. Methods such as behavioral cloning [204, 251] minimize the discrepancy between predicted and expert actions, while inverse reinforcement learning [252, 253] aims to infer the underlying reward function from demonstrations. Although IL accelerates learning by reducing exploration, it struggles with covariate shift, where small errors during policy execution compound over time, which limits generalization in novel scenarios [254].
A hybrid approach that combines RL and IL leverages the strengths of both. Typically, policies are initialized with expert demonstrations to accelerate training using IL, and then refined by RL to adapt to dynamic environments [255, 254, 256]. Additionally, methods such as Guided Policy Search [233] or Reward Shaping [234] integrate demonstration data into the RL process, enhancing both data efficiency and generalization. These combinations bridge the gap between the efficiency of the IL and the adaptability of the RL, improving the low-level actuation capability to execute actions in real-world scenarios.
@@ 2.3 High-level Planning
Low-level actuation can only meet the needs of simple, reactive tasks, but it struggles to handle the complexity of planning long-horizon tasks with multiple sub-tasks. To address this limitation, high-level planning algorithms have been developed. Classic planning algorithms, such as A* algorithm [257, 258, 259, 260], Dijkstra’s algorithm [261, 262, 263], and the probabilistic roadmap (PRM) approaches [264, 265, 266, 267], serve as foundational methods in this domain. Despite their significant influence and effectiveness in structured environments, these algorithms encounter substantial challenges in real-world scenarios, particularly in high-dimensional state spaces and under conditions of partial observability. Recent research has explored the use of LLMs as high-level planners in embodied systems, bridging cognitive reasoning and physical task execution by translating abstract instructions into actionable robotic tasks [268, 64]. Table 2 summarizes representative approaches of different high-level planning methods.
@@
LLM-based task planners typically break down high-level goals into a sequence of executable subtasks [269, 270, 271]. There are generally two main paradigms: code-based planners and language-based planners, as shown in the Figure 4. Code-based planners [272, 273, 274] operate by selecting from a pre-defined set of modular skills or functions, invoking them via APIs to execute tasks step by step. They excel in highly controlled workflows requiring safety and reliability, where predefined APIs or modules can ensure predictable outcomes. Representative code-based planners such as the DEPS framework [275] further emphasize “self-explanation” to better exploit the capabilities of LLMs, where an LLM generates plans, explains failures, and uses environmental feedback to aid in re-planning. In addition, by transforming observation sequences into 3D scene graphs using visual language models, ConceptGraphs [274] helps the LLM reason about spatial and semantic relationships in task planning. These methods benefit from the deterministic nature of pre-programmed functions, reducing ambiguity in execution. However, their dependency on predefined skills makes them less adaptable to unexpected changes or tasks outside the pre-programmed domain.
@@
Language-based planners [276, 277, 278] show greater flexibility without using predefined functions. In open-ended domains like creative problem-solving, customer support, or planning in uncertain environments, language-based planners can use contextual cues to refine and adjust their instructions in real-time. Additionally, these methods dynamically integrate diverse feedback—ranging from success indicators to human inputs—without necessitating additional training for the LLMs. Such adaptive feedback and re-planning mechanisms [279, 275, 280] enable systems to recover from unexpected states, offering more flexibility to novel environments. This adaptability, however, can sometimes come at the cost of consistency or precision [123]. Moreover, by integrating MLLMs’ capabilities, EmAI systems like Socratic models [281] effectively translate non-language inputs into unified language descriptions through multimodal informed prompting. This approach not only streamlines information exchange across different modalities through MLLMs but also enhances robot perception and planning tasks [282, 123, 283].
@@ 2.5 Synergistic Integration
The aforementioned four key functionalities are often initially developed independently but must be effectively integrated to construct a comprehensive EmAI system. Integration approaches such as MemoRAG [336] and Reflexion [337] enhance high-level planning processes by retrieving relevant information [338, 339, 340] or summarizing past experiences [341, 342] from memory modules. These methods improve adaptability to dynamic environments and enable more rational decision-making. Closed-loop approaches, including RoboGolf [343], LyRN [344], and AlphaBlock [345], integrate perception and actuation modules, leveraging feedback to refine observations and dynamically update control signals. This integration facilitates precise action adjustments and supports effective multi-step planning. Moreover, active and interactive perception systems [185, 43, 346] take this a step further by engaging in real-time interactions to explore object properties, update environmental contexts, and refine decisions based on immediate outcomes. The resulting actions and observations are stored in the EmAI memory, which can be used to construct multimodal knowledge graphs integrating physical properties, concepts, affordances, and intentions for future use [347]. The modern AI alignment approaches that combine modules [348, 349] and foundational models that encompass all functionalities of these modules [350, 351] are also being widely researched as promising areas [352, 353]. However, there is still a lack of a highly compatible, efficient, and effective unified architecture capable of integrating various developed modules. Achieving alignment and seamless integration among these modules, while minimizing development (e.g., fine-tuning), remains an open challenge.
@@ 3.1.1 Pre-Intervention Stage
Virtual triage nurse. In modern smart healthcare systems, EmAI-based virtual triage nurses are replacing human nurses and play a pivotal role in streamlining patient management by directing individuals to the most appropriate clinical departments. These EmAI systems analyze patient-reported symptoms and rank departments based on symptom descriptions [359, 360, 361]. Beyond symptom-based sorting, advanced triage systems integrate wearable health data and EHR to provide a holistic assessment of conditions [362, 363, 364]. In emergency care settings, these systems can even predict patient outcomes and recommend intervention pathways, significantly shortening response times during critical situations [365, 366, 367, 368]. As healthcare systems become more interconnected, intelligent triage systems are increasingly serving as an efficient tool for future infrastructure [364, 369], enabling seamless coordination between primary care, specialist consultations, and hospital admissions.
@@
Interactive medical consultant. Recently, some LLM-based chatbots, such as DISC-MedLLM [370] and HealAI [371], have been developed to provide instant, reliable, and context-specific responses to medical inquiries, helping patients better understand their symptoms, treatment options, or follow-up care [372, 372, 373]. They can also explain medical conditions, offer personalized recommendations [374, 375], and explain radiology reports [376]. By bridging the gap between patients and complex medical knowledge, interactive medical chatbots not only empower individuals to make informed decisions but also reduce the workload of healthcare professionals [377, 378]. With advancements in LLM reasoning [379], these systems are becoming an essential component of patient-centered smart healthcare solutions. Beyond answering medical questions, these chatbots can also guide patients through administrative processes, such as booking appointments [380, 381] or managing prescriptions [382, 383], thereby easing the burden on healthcare staff [384, 385]. With their ability to operate around the clock, these systems foster greater accessibility and trust in the medical process.
@@
Imaging analyst. Another representative area is robot-assisted medical imaging, which not only improves the accuracy of diagnostic processes but also expands the capabilities of medical imaging in complex anatomical assessments. Many studies have focused on medical image analysis using AI techniques [386, 387, 388], but to support robotic surgery and preoperative robotic diagnosis requires additional requirements including real-time processing [389], 3D spatial understanding [390], and safety compliance [391] beyond high precision. Among medical imaging technologies including ultrasonography, radiology, and endoscopy, EmAI enhances their capabilities in different ways.
@@
For ultrasonography, which is portable, real-time, non-invasive, and relying on the synchronization of diagnostic and procedural operations, EmAI can serve as a remote assistant to help clinicians conduct remote ultrasound diagnosis [105] and protect themselves from risks against epidemics [392]. For radiology, EmAI can also help clinicians in lesion localization [77, 393, 394], surgical planning [395, 396, 397], intraoperative navigation [398, 399] with a better understanding of the raw imaging and attached report from X-Ray, Computed Tomography (CT), and Magnetic Resonance Imaging (MRI). For endoscopy, EmAI can process endoscopic video feeds in real-time to identify abnormalities [400, 401], such as polyps or early-stage tumors, reduce operator fatigue, enable precise navigation via path optimization [402, 403], and assist in procedural tasks like polypectomy with adaptive, precision-controlled robotic movements [404, 405].
@@
    Endoscopic navigator. The integration of EmAI in endoscopic procedures has significantly enhanced the precision and efficiency of minimally invasive diagnostics and treatments by providing real-time guidance for device positioning and lesion targeting [426, 427]. For example, AI-powered 3D mapping techniques reconstruct anatomical structures in real time, enabling clinicians to navigate complex regions such as the gastrointestinal tract and respiratory pathways with greater accuracy [428]. Robotic-assisted endoscopy further utilizes automated control systems that adapt dynamically to patient-specific anatomy and procedural demands [429, 430], which improves precision in tasks such as biopsy, polyp removal, and other targeted interventions. Furthermore, these systems incorporate predictive analytics to refine surgical pathways, reduce tissue trauma, and streamline procedural workflows, enhancing safety and operator confidence in endoscopic interventions [431, 432].
In summary, fully autonomous ultrasound system [110], human-centric radiology assistant [433], and flexible robotic endoscope system [434] capable of navigating and adjusting diagnostic examinations with minimal human supervision have been developed, achieving a higher level of intelligence and autonomy. Compared to human-led examinations, these systems optimize workflows, improve patient care quality, and ensure a safer environment for all involved in the surgical process. Additionally, these systems can also function as a module within a more comprehensive EmAI system such as surgical robots, showcasing opportunities to expand their applications.